### v1

# 1Manifold Research 2Metarch.ai 3Georgia Tech

Abstract
Vision-language-action (VLA) models represent an important step toward general-purpose robotic systems by integrating visual perception, language understanding, and action execution. However, systematic evaluation of these models, particularly their zero-shot generalization capabilities in procedurally generated, out-of-distribution (OOD) environments, remains limited. In this paper, we introduce MultiNet v0.2, a comprehensive benchmark designed to evaluate and analyze the generalization performance of state-of-the-art VLM and VLA models—including GPT-4o, GPT-4.1, OpenVLA, Pi0 Base, and Pi0 FAST—on diverse procedural tasks from the Procgen benchmark. Our analysis reveals several critical insights: (1) all evaluated models exhibit significant limitations in zero-shot generalization to OOD tasks, with performance heavily influenced by factors such as action representation and task complexity; (2) OpenVLA generally outperforms other models due to its robust architectural design; and (3) GPT variants demonstrate substantial improvements when constrained appropriately, highlighting the sensitivity of model performance to precise prompt engineering.
I Introduction
Recent advancements in large-scale vision-language models (VLMs) and vision-language-action models (VLAs) have demonstrated remarkable capabilities across various domains, including image recognition, natural language understanding, multimodal association, and preliminary robotics applications [59, 39, 10]. These models promise a future where general-purpose AI systems can interpret visual inputs, comprehend language commands, and execute appropriate actions in diverse scenarios.
However, a significant challenge remains: ensuring these models can generalize effectively to out-of-distribution (OOD) tasks. Current models often struggle with zero-shot transfer capabilities, particularly when confronted with novel scenarios or tasks that differ substantially from their training data [31]. This limitation is especially evident in procedurally generated environments, such as those provided by the Procgen benchmark[14], which are designed to test visual understanding, decision-making, and action generation capabilities in varied and unpredictable settings.
The architecture and training methodologies of these models play a crucial role in their generalization abilities. Factors such as the nature of the action space (continuous vs. discrete), the domain of training data (real-world robotics vs. simulated environments), and the methods of input-output processing can significantly impact performance [30]. For instance, models trained predominantly on robotics data may not perform well in simulated game environments due to differences in action representation and environmental complexity.
To address these challenges, we introduce MultiNet v0.2, a comprehensive benchmarking effort aimed at evaluating the generalist capabilities of VLMs and VLAs on procedurally generated tasks. Our study encompasses a diverse set of models, including GPT-4o, GPT-4.1, OpenVLA, Pi0 Base, and Pi0 FAST, assessed across multiple Procgen datasets[52, 53, 39, 6, 37, 14].
Our contributions are threefold:
  • We provide a systematic benchmarking framework that evaluates the performance of state-of-the-art models on a variety of procedurally generated environments.
  • We analyze the impact of architectural choices ,training data, and output processing techniques on model generalization, highlighting the limitations of current approaches.
  • We offer insights into how factors like action space representation and image complexity influence model performance, paving the way for future improvements in generalist AI systems.
The remainder of this paper is structured as follows: Section 2 reviews related work in the field; Section 3 details our experimental setup and datasets; Section 4 presents our empirical results and analysis; and Section 5 discusses the implications of our findings and outlines directions for future research.
II Related Work
Our work builds on the rapid advancement of physical and digital vision-language-action models that enable agents to interpret, reason, and act across diverse tasks and embodiments. In this section, we review the landscape of related work, organized into four key areas with an emphasis on digital environments: existing action models, datasets and benchmarks, data collection tools, and agentic applications. For comprehensive surveys on VLAs and computer use agents, see [36, 45, 49, 61].
Action models. Vision-language-action (VLA) architectures generally consist of four core components: vision encoders, language encoders, planners, and action decoders. Recent work has advanced each element to improve end-to-end performance.
First, on the encoder and fusion side, richer multimodal representations have been achieved by integrating advanced vision backbones and efficient conditioning mechanisms. For example, SigLIP [82] image embeddings have been adopted in OpenVLA [39] and PaLiGemma [4], while FiLM [55] conditioning enables RT-1 [9], BC-Z [38], and OpenVLA-OFT [40] to modulate visual features with language cues. More recently, Interleave-VLA [21] interleaves vision and language embeddings at multiple layers to improve cross-modal alignment.
Second, the planner, sitting between perception and action, can be categorized into three paradigms.
  • World model planner, which first learns a compact dynamics model in a latent space and then ’imagines’ trajectories for policy extraction. Representative examples include Dreamer series [32], Genie series [20], DINO-WM [85], AdaWorld [25], and Latent Diffusion Planning (LDP) [75].
  • Sequence-based action generation, where control is treated as conditional sequence prediction by large transformers. This category unifies (a) pretrained multimodal decoders, such as RT-2 and OpenVLA, which generate actions autoregressively over vision–language embeddings [10, 39], and diffusion-policy methods like Octo [26], TinyVLA [72], $\pi_{0}$ [6], CogACT [44], and GR00T N1 [5], with (b) pure sequence planners like the Decision Transformer and its hierarchical variants, which autoregressively predict future actions from past state–action histories and a return-to-go token [13, 15].
  • Reactive model-free control, in which policies are learned directly from demonstrations or offline data without any explicit planning module. Classic methods include Behavior Cloning[56], DAgger [60], and GAIL [33] and new variants such as Implicit Behavior Cloning[23], which excel in short-horizon or well-covered tasks but lack the capability for roll-out–based foresight.
Each paradigm balances sample efficiency, generalization capability, and the degree of explicit world modeling: model-based planners offer strong long-horizon reasoning at the cost of model learning effort; sequence-based methods leverage powerful pretrained sequence models to bypass explicit dynamics but depend on large-scale pretraining; and model-free approaches provide simplicity and stability in familiar settings but struggle with unobserved or out-of-distribution scenarios.
Specifically, to handle long-horizon and hierarchical tasks, many VLA systems combine:
  • Speed–accuracy cascades: two-model pipelines with fast/slow predictors[6, 7, 68, 22, 5].
  • Planning precision: chain-of-thought reasoning with VLAs[84].
  • Horizon compression: hierarchical vector quantization in VQ-BeT[42].
  • Action-chunking: parallel decoding in OpenVLA-OFT and FAST token compression[39, 37].
  • Latent planning: AgiBot GO-1 ViLLA framework with a VQ-VAE latent model and a 24-layer transformer[68].
These architectural advances have enabled VLAs to excel not only in standard benchmarks but also across very different application domains. In robotics, OpenVLA [39] is fine-tuned to excel in 29 multi-object manipulation tasks (e.g., block stacking, pick-and-place, table cleaning); VQ-BeT [42] performs multistep kitchen behaviors such as drawer opening/closing and toaster/fridge operation; and the $\pi$-series [6, 37, 7] demonstrate zero-shot generalization on tasks like bowl stacking, towel folding, and grocery bagging. In digital environments, WebGUM [24], OpenAI Operator[54], Claude Computer Use [2], Manus AI [67], and CogAgent [35] can navigate GUIs and carry out web tasks such as clicking, typing, filling in forms, searching, and booking, while Ferret-UI [81] excels at low-level interface understanding (e.g., icon and widget detection). Notably, unified models like Magma [79] that bridge digital GUI actions with physical-world manipulation still remain underexplored.
Datasets & benchmarks. Recent embodied VLA work utilizes large-scale, multitask robot datasets and benchmarks such as OpenX [51], LIBERO [48], BridgeData [18, 19], VLABench [83], Agibot World [68], egocentric benchmarks such as EGO-4D [28], augmented datasets with affordance, action trajectory, and object annotations [16, 79], and video-action datasets such as Something-SomethingV2 [27], Motion-X [46], and MotionBank [77] using actions annotations or relying on architectural components [43, 80] to learn actions from human demonstration videos. In the domain of OS and browser use, GUI agents are evaluated on web pages, desktop UIs with datasets such as Mind2Web (Web navigation tasks) [17], AITW (Android device control) [58], WebUI (400 K rendered pages + metadata) [73], OSWorld (real-computer environment tasks) [76], GUI-World (6 GUI scenarios, dynamic content) [12], and VideoGUI (instructional‐video–driven GUI editing) [47].
Data collection tools. Compared to the abundant text training data on the Internet for natural language models, the lack of action datasets [41] requires proper data collection tools for VLAs. In robotics, some mobile teleoperation interfaces were engineered for tasks such as 6-DoF demonstration [50], and multi-arm coordination [70]. As for computer use agents, commercial RPA tools such as UiPath and Automation 360 capture user interactions with applications, including clicking, reading, and writing actions on user interface objects while open-source browsers (Selenium IDE, Playwright Codegen, Browserless Recorder) capture browser operations. On the OS level, open source tools such as selfspy [63], can capture keystrokes, visited URLs, and executed commands. Even though these emerging tools extend our ability to collect large-scale computer-use action datasets, the tools for action data collection in digital VLAs are still lacking, and building large-scale, high-quality action datasets remains a major challenge.
Agentic applications. Digital agentic systems now operate across desktop and mobile platforms as well as within web browsers, serving domains from customer support, finances, to healthcare and enterprise productivity.
Closed source, generalist multimodal computer use agents such as OpenAI Operator, Claude Computer Use, Manus AI interact with GUIs and browser content using only image and text inputs to automate workflows end-to-end[54, 2, 67]. Open source agents such as AutoGPT [3] and OS-Copilot [74] interface with terminals, file systems, and third-party applications, while specialized web agents [24, 29] focus on browser navigation and form automation.
In enterprise contexts, CRM and support workflows are increasingly driven by agents like Salesforce Einstein GPT [62] and ServiceNow Virtual Agents [65] ground responses in enterprise data to automate case routing, conversation summarization, and triage escalations. In healthcare, tools such as Google’s Articulate Medical Intelligence Explorer (AMIE) [69] demonstrate text-based diagnostic performance on par with or exceeding primary care physicians in controlled studies.
Despite these advances, current LLM-based agentic applications still struggle with reliability, accuracy, and efficiency when faced with dynamic, complex, and uncontrolled real-world environments due to their fundamentally incomplete world model [71], tendency toward self-deception [78], and lack of appropriate bounds on agency [11].
III Experimental Setup
III-A Dataset
We utilized offline trajectories from expert reinforcement learning (RL) agents trained on the Procgen dataset, which were sourced from Facebook’s publicly available repository  [^1] https://dl.fbaipublicfiles.com/DGRL/1M/expert/ [14]. To facilitate consistent and straightforward data handling, we translated these trajectories into the TensorFlow Datasets (TFDS) format [^2] https://www.tensorflow.org/api_docs/python/tf/data/Dataset. This standardized format ensured streamlined loading and usage across all subsequent profiling experiments.
For evaluation purposes, we established test splits comprising 10 percent of the episodes randomly sampled from each subdataset. Given that the Procgen dataset consists of 16 distinct subdatasets, this approach resulted in balanced and representative test splits for comprehensive evaluation. Importantly, all profiling and performance assessment experiments reported in this study were conducted exclusively on these designated test splits, ensuring that our results reflect the models’ genuine generalization capabilities without contamination from training data.
III-B Models
In this study, we leveraged state-of-the-art Vision-Language-Action (VLA) models with predefined full-sized weights to evaluate performance comprehensively. The models and their respective sources were:
OpenVLA: The weights were obtained from the HuggingFace repository under the identifier openvla/openvla-7b.
Pi0 Base: We utilized weights stored at s3://openpi-assets/checkpoints/pi0_base/params. We obtained this path from the Openpi open-source codebase [^3] https://github.com/Physical-Intelligence/openpi
Pi0 Fast: We used weights accessible via s3://openpi-assets/checkpoints/pi0_fast_base/params. We obtained this path from the Openpi open-source codebase.
All models were employed in their complete forms, without any quantization techniques, to ensure accurate and undiluted performance evaluation.
To align the models with the Procgen task context and dataset specifics, we performed targeted adaptations during ingestion and inference.
III-B1 Data Ingestion
GPT 4o and GPT 4.1: Employed the Genesis prompt engineering framework [^4] https://github.com/ManifoldRG/MultiNet/tree/main/src/modules, meticulously crafting timestep-specific prompts that including detailed instructions and constraints for inference, explicit descriptions of the subdataset task and environment, visual observations corresponding to the current timestep, clear specifications of the action space and range, and verbal descriptions for each possible action. As seen in the Appendix section VII-C the prompt constructed for GPT 4.1 was improved over the v0.1 Genesis prompt to observe the performance increase of VLMs due to a more rigorously curated prompt. The information given to the VLMs was the same through the old and new prompts, but the new one includes changes to ensure more valid outputs from the model. The new prompt explicitly includes the fact that this is a simulation dataset. The new prompt also uses capitalization to emphasize instructions. Additionally, rather than stating all the possible types of inputs that could be given from different types of action datasets in Multinet, the new prompt includes only the type of inputs that will be given as a part of the Procgen datasets.
OpenVLA: Received visual observations per timestep as direct image inputs, accompanied by succinct single-line text prompts describing the corresponding subdataset task.
Pi0: Incorporated the first image view as the primary visual observation, supplemented with zero arrays for the second and third image views, with image masks indicating True for the primary view and False for others. Proprioceptive states were also set as zero arrays. Each timestep included a concise single-line task description prompt.
Pi0 with Fast: Similar to Pi0 Base, Pi0 Fast utilized the first image view as its primary visual observation but employed zero arrays for the second and third image views with image masks set to True for all three views. Zero arrays represented the proprioceptive states, and each timestep included a single-line task description prompt.
III-B2 Model Adaptation
OpenVLA: Adaptation involved restricting the autoregressive step to one, ensuring single-dimensional action vector predictions. Generated actions were unnormalized using statistics computed across the entire Procgen subdataset and subsequently rounded to discrete actions. For Brier MAE calculation, logits from the Llama 2 backbone were extracted, probabilities computed, and token mappings from the Llama vocabulary to Procgen integer actions were created and cached to facilitate efficient inference.
Pi0 Base: The action horizon was constrained to one timestep, with flow matching denoising executed over ten steps (default configuration). Actions were predicted with a default dimension of 32. The first dimension of the predicted action was picked to ensure a final single dimensional action vector prediction, which was then unnormalized based on Procgen dataset-wide statistics (same dataset statistics used for OpenVLA), and discretized by rounding.
Pi0 Fast: Adaptation entailed adjusting the action horizon to one timestep and setting the action dimension to one, compatible with Procgen’s single-dimensional action space. The decoding steps were limited to four tokens (”Action”, ”:”, a space character, and the Paligemma (VLM backbone of Pi0 models) location token) to explicitly generate the correct action value. To compute probabilities for Brier MAE, leveraging to the static nature of the BPE tokenizer, token mappings between the Paligemma token IDs and Procgen integer actions were established post-training. During inference, probabilities extracted from the Paligemma backbone logits were aggregated according to this mapping. Recognizing Pi0 Fast’s inference speed was significantly slower (approximately ten times slower than OpenVLA), we optimized runtime by caching embeddings for the two static zero-images after initial processing by the SigLIP embedding, effectively doubling inference speed.
III-C Inference Infrastructure
For systematic performance evaluation and model profiling, we utilized dedicated hardware resources optimized for each model’s computational requirements:
  • OpenVLA inference was executed on an NVIDIA L4 GPU instance, selected for its balance of computational efficiency and memory capacity, ideal for models of intermediate complexity.
  • Pi0 Base utilized a single NVIDIA A100 GPU instance equipped with 40 GB of memory, providing ample computational power and memory bandwidth for accurate and efficient inference.
  • Pi0 Fast, due to its greater computational demands and inference complexity, was allocated four NVIDIA A100 GPU instances, each with 40 GB of memory, facilitating parallel processing to substantially enhance inference throughput.
  • In contrast, GPT 4x inference was managed externally through batch job APIs, eliminating local processing overhead and leveraging scalable cloud resources to efficiently handle its computational requirements.
III-D Evaluation metrics
In order to capture the ability of state-of-the-art VLAs and VLMs to generalize to completely unseen, OOD action data, we carefully curated a list of metrics to evaluate the models as fairly as possible. These metrics capture a variety of aspects of performance of the models such as how well calibrated they are in their estimates, reliance on outliers, how they get affected by class imbalance, how well they are able to restrict their predictions to the valid action space, whether the models are biased towards certain kinds of subdatasets or actions, etc.
  • Brier Mean Absolute Error
    Brier $\text{Brier score}=\frac{1}{N}\sum_{t=1}^{N}\sum_{i=1}^{R}|f_{ti}-o_{ti}|$ (1)
    where $f_{ti}$ is the probability of the model prediction for class i at timestep t, $o_{ti}$ is the $i^{th}$ dimension of the one hot representation of the ground truth action for the same timestep, R is the possible number of classes, and N is the total number of timesteps.
    Brier MAE is a variation of the original Brier score [8], which is a useful method to measure the accuracy of probabilistic predictions. In our case, the probability that the action in a given timestep is one out of all the possible actions in the action space. It gives a good idea of how well the predictions of the models are calibrated. Brier score penalizes predictions that are both wrong and confident more severely than those that are wrong but uncertain. Additionally, unlike Mean Squared Error, using Mean Absolute Error maintains the original scale of the data making it potentially more interpretable in the action space’s context. MAE is also less sensitive to outlier predictions than squared error metrics, thus giving a more stable evaluation of typical performance. Maximum possible Brier score is 2, which is assigned to invalid predictions as a high penalty.
  • Normalized Brier Mean Absolute Error Average of Brier Absolute Errors that have been min-max normalized using the minimum and maximum Brier absolute errors over all timesteps. A discrepancy in trends between regular Brier MAE and Normalized Brier MAE indicates reliance on outlier predictions that lead to abnormally high or low Brier MAEs.
  • Normalized Quantile Filtered Brier Mean Absolute Error This metric value is obtained by filtering out the Brier Absolute Errors that are greater than or equal to the 5th percentile error or lesser than or equal to the 95th percentile error, and then normalizing based on the quantile filtered minimum and maximum errors. This metric helps indicate how good or bad the majority of the predictions are. Discrepancy in trends between normalized and normalized quantile filtered Brier MAEs indicates the reliance on a few extremely good or extremely bad predictions.
  • Maximum relative Brier Mean Absolute Error
    Max relative Brier $\text{Max relative Brier MAE}=\frac{\max(MAE)}{\text{median}(MAE)}$ (2)
    This metric quantifies how much the worst case error deviates from the typical (median) error produced by the model for a given subdataset of Procgen. A value significantly greater than 1 indicates the presence of some extremely bad predictions that can skew the results.
  • Micro Precision, Recall, F1 score, Exact Match rate
    Micro Precision=True PositivesTrue Positives + False $\text{Micro Precision}=\frac{\text{True Positives}}{\text{True Positives + False Positives}}$ (3)
    Micro Recall=True PositivesTrue Positives + False $\text{Micro Recall}=\frac{\text{True Positives}}{\text{True Positives + False Negatives}}$ (4)
[…]
    Exact Match Rate=Num timesteps where preds = labelsTotal number of $\text{Exact Match Rate}=\frac{\text{Num timesteps where preds = labels}}{\text{Total number of timesteps}}$ (6)
    For our case, where it is a multi-class, single correct prediction task and the classes (actions) are mutually exclusive, false positives are equal to false negatives. This is because an incorrect predicted action in a given timestep is a false positive for its class and a false negative for the ground truth action class. Due to this, precision = recall = f1 score = exact match rate when calculating micro metrics in our case. The micro precision/recall/f1 is calculated over all timesteps of the test split of a given subdataset. This metric measures how precise the models are - high precision indicates that the model gets a high number of predictions correct, and a low precision suggests the model is frequently predicting actions that don’t match the ground truth.
  • Percentage invalids Predictions that fall outside the valid action space of a given subdataset of Procgen are considered invalids. In the case of the VLMs - GPT 4o, and GPT 4.1, the outputs that are not in the desired format of an action space size vector containing discrete values as its elements, are also considered invalids. Thus, percentage invalids is calculated as
    Percentage invalids=Timesteps with invalid predsTotal number of $\text{Percentage invalids}=\frac{\text{Timesteps with invalid preds}}{\text{Total number of timesteps}}*100$ (7)
    A high invalid percentage indicates that the model struggles to produce outputs in the target action space.
  • Micro Precision without Invalids This metric is a variation of micro precision where the false positives are only counted for valid predictions. By not considering the invalid predictions, this metric gives an idea about the model’s performance in the cases where it is able to produce valid outputs. A discrepancy in the trends between micro precision and micro precision without invalids indicates that a model would be more capable if it was fine-tuned, prompted, or had its output processed to constrain the predictions to the valid action space.
  • Class-wise Precision, Recall, F1 scores Class-wise variations of the precision, recall, and F1 metrics are the same metrics calculated for each action class within the valid action space of a subdataset individually. The class-wise precision, recall, and F1 scores of a model would be calculated for a given action class x subdataset combination. Average class-wise metrics for a model are calculated by averaging the class-wise x subdataset metric values over all subdatasets to report a unified set of class-wise metrics for a given model.
    These metrics help understand model biases and preferences towards specific action classes.
  • Macro Precision, Recall, and F1 scores
    Macro precision, recall, and F1 metrics are calculated by averaging the class-wise metrics across all classes in the valid action space of a given subdataset for a given model. These metrics are very important as they help alleviate the effects of the dominating majority class that skews performance. With micro metrics, the majority classes are favored due to global aggregation. This leads to a misunderstood notion of performance, especially in the Procgen expert data, as there are numerous cases of class imbalance which can in turn skew the values of the micro metrics. Macro metrics prove to be a better indicator of a model’s performance when compared to micro metrics irrespective of the nature of the test dataset as they treat all classes equally regardless of size. Within this, Macro recall is typically the most representative metric for performance as it is neither affected by majority classes as in the case of micro metrics, nor is it affected by the performance of models on the rare classes as in the case of macro precision or F1 scores.
IV Results and Discussion
IV-A Models Exhibit Poor Zero-Shot Out-of-Distribution (OOD) Generalization
IV-A1 Comparing General Performance
In evaluating the generalization capability of GPT 4o, GPT 4.1, OpenVLA, Pi0 Base, and Pi0 FAST, we observed universally poor performance on zero-shot out-of-distribution (OOD) tasks. The analysis using various performance metrics underscored significant limitations across all models.
As seen in Figure 1, The Brier Mean Absolute Error (MAE) scores revealed poor probability calibration, as scores approached the maximum value of 2 across all models and datasets. Specifically, GPT 4o, GPT 4.1, and Pi0 FAST consistently scored above 1.70, while OpenVLA generally exceeded 1.50. The subdatasets on which models exhibited the weakest performance were Starpilot for GPT 4o, Leaper for GPT 4.1, Miner for Pi0 FAST, and Ninja for OpenVLA.
Figure 1: Brier Mean Absolute Error scores across 4 models - GPT 4o, OpenVLA, GPT 4.1, and Pi0 FAST. Pi0 Base is a diffusion based model and can not be evaluated using Brier MAE due to lack of logits in its inference architecture. All models display Brier MAE close to 2 indicating poor performance.
Analyzing micro-level metrics, precision values were low across all models and subdatasets. OpenVLA achieved the highest micro precision at approximately 25 percent on the Coinrun dataset, whereas GPT 4o showed the lowest precision at 0.9 percent on Starpilot. Pi0 FAST reached a minimal precision of around 1 percent on Heist. Overall, as can be inferred from Figure 2, the performance hierarchy based on micro precision was: OpenVLA $>$ Pi0 Base $>$ GPT 4.1 $>$ GPT 4o $>$ Pi0 FAST.
Figure 2: Micro precision across all 5 models - GPT 4o, OpenVLA, Pi0 Base, Pi0 FAST, and GPT 4.1. When considering all 16 subdatasets, Pi0 FAST overall showcases the worst performance, and OpenVLA the best performance.
When considering micro precision without invalid predictions, as seen in Figure 3 significant improvements were noted for GPT 4o and Pi0 FAST, highlighting their potential if constrained or further fine-tuned to produce valid action outputs. GPT 4o’s precision notably improved, surpassing Pi0 Base in a few subdatasets such as Coinrun, Bossfight, and Dodgeball. Nevertheless, the adjusted performance order still positioned OpenVLA at the top, followed by Pi0 Base, Pi0 FAST, GPT 4o, and GPT 4.1.
Figure 3: Micro precision without considering invalids across all 5 models. GPT 4o and Pi0 FAST show a significant surge in precision once invalids are not considered.
Macro precision analysis revealed that GPT 4o, OpenVLA, and Pi0 Base maintained similar performance levels across most datasets. As seen in Figure 4, conversely, GPT 4.1 and Pi0 FAST demonstrated significant variability, implying distinct preferences for specific task types. Despite these observations, macro precision remained generally low across all models, with a maximum slightly above 15 percent. Importantly, comparing its macro and micro precision performances, it is clear that OpenVLA’s performance heavily favored majority classes, while GPT 4o exhibited relatively stronger performance on minority classes.
Figure 4: Macro precision across all 5 models. GPT 4.1 and Pi0 FAST show variability and preference towards specific subdatasets. OpenVLA shows reliability on majority classes whereas GPT 4o displays relatively stronger performance on minority classes.
Macro recall, as seen in Figure 5 highlighted OpenVLA’s aggressive prediction strategy, resulting in consistently high recall but lower precision due to numerous false positives. GPT 4o, despite higher macro precision, showed substantially lower recall indicating pronounced class biases towards a few classes, and a potentially overly conservative nature about predicting certain classes compared to other classes. GPT 4.1 exhibited a relatively balanced approach, achieving both substantial recall and precision, thus indicating less bias towards specific classes. Pi0 Base displayed a moderately balanced performance without leading in any subdataset, whereas Pi0 FAST consistently showed low recall, reflecting challenges in identifying true positives across all classes.
Figure 5: Macro recall across all 5 models. OpenVLA performs better when considering macro recall compared to macro precision, indicating a high number of false positives. GPT 4o shows lower macro recall than precision indicating biased performance towards specific minority classes. GPT 4.1 and Pi0 Base show relatively less biased and moderate performance, whereas Pi0 FAST showed consistent low recall.
The analysis of invalid predictions highlighted significant disparities. OpenVLA, by design, produced zero invalid predictions, ensuring actions always fell within the valid action space. As seen in Figure 6 GPT 4o and Pi0 FAST frequently generated invalid outputs, with some datasets reaching invalid prediction rates above 80 percent. Conversely, GPT 4.1 and Pi0 Base demonstrated notably lower invalid prediction rates, underscoring their relatively stronger capability to adhere to the valid action space constraints.
Figure 6: Percentage invalids across all 5 models. Invalids refer to model predictions that are not valid actions in the subdataset’s action space. Pi0 FAST and GPT 4o struggle to produce valid actions irrespective of the subdataset
IV-A2 Outlier Influence and Majority Prediction Performance
As seen in Figures 7 and 8 the normalized Brier MAE and quantile-filtered analysis indicated consistent error trends when compared to regular Brier MAE scores across models and datasets, demonstrating minimal reliance on outlier predictions. Given the already high median errors, the influence of outliers was negligible.
Figure 7: Normalized Brier MAE across 4 models - GPT 4o, OpenVLA, GPT 4.1, Pi0 FAST. Follows trends of regular Brier MAE indicating minimal reliance on outliers
Figure 8: Normalized Quantile Filtered Brier MAE across 4 models. Follows trends of regular and Normalized Brier MAE scores indicating minimal reliance on extreme predictions
Further investigation using the max relative Brier MAE metric showed minimal variance in model performance, the scores staying consistently close to 1.0. This finding, seen in Figure 9, suggests limited deviation between median and extreme error values, primarily due to the universally high error levels across all models, which reduce the relative significance of outlier predictions.
Figure 9: Maximum relative Brier MAE across 4 models. All values are equal to or around 1.0, indicating limited deviation between median and extreme bad predictions.
IV-B Models Struggle with Unseen Discrete Action Data
IV-B1 Dataset-Specific Performance Across Models
Our evaluation revealed distinct performance patterns across the five models—OpenVLA, GPT-4o, GPT-4.1, Pi0 Base, and Pi0 FAST—when exposed to out-of-distribution (OOD) datasets with discrete action spaces from Procgen.
OpenVLA consistently achieved the highest macro-recall among the evaluated models across nearly all 16 Procgen subdatasets, typically ranging from 9 percent to 12 percent. It performed notably better on simpler tasks, achieving a maximum recall of approximately 12.5 percent on the Maze dataset, characterized by straightforward navigation and collection actions without the complexity of additional special actions. Conversely, OpenVLA’s recall dropped to its lowest ( 8 percent) on the Ninja dataset, which involves multiple directional special ”fire” actions, indicating difficulty handling tasks requiring timely execution of specialized actions.
GPT-4.1 and Pi0 Base occupied a middle performance tier, displaying macro-recall values roughly between 6 percent and 11 percent. GPT-4.1 generally outperformed GPT-4o across most datasets and showed superior performance compared to Pi0 Base on simpler datasets like Leaper, Climber, Jumper, Maze and Heist. Conversely, GPT-4o exhibited significantly weaker performance than Pi0 Base on datasets involving timely and sparse execution of special actions such as Plunder, Fruitbot, and Starpilot. These three datasets involve ”fire” actions, indicating GPT-4o’s difficulty in handling timely execution tasks. Specifically, GPT-4.1 achieved macro-recall scores around 11 percent on simpler datasets including Leaper, Climber, Jumper, Bigfish, Chaser, Coinrun, Miner, Maze, and Heist. In contrast, its performance diminished to around 7 percent on more challenging environments such as Plunder, Starpilot, Bossfight, and Ninja. GPT-4o showed the highest recalls around 8.5 percent on datasets with simpler movement-only actions (Leaper, Climber, Jumper, Coinrun, Maze, Heist), but performed poorly (around 3 percent) on more complex tasks like Chaser and Bossfight, and extremely poorly (below 2.5 percent) on Fruitbot and Starpilot, both featuring special actions.
Pi0 Base achieved macro recalls typically between 6 percent and 10 percent. Its best performance (around 10 percent) appeared on datasets such as Leaper, Plunder, Jumper, Bigfish, Chaser, Coinrun, and Miner. Conversely, macro recall dropped to approximately 6 percent on datasets involving special actions like Dodgeball, Fruitbot, Bossfight, and Ninja.
Pi0 FAST exhibited the lowest macro recall among the evaluated models, frequently below 6 percent, and occasionally as low as 2 percent on datasets including Leaper, Bigfish, Coinrun, Ninja, Miner, Maze, and Heist. However, Pi0 FAST displayed anomalously higher recalls ( 10 percent) on datasets such as Starpilot and Bossfight. Despite these datasets involving special actions, the actions are executed frequently rather than sparsely, aligning better with Pi0 FAST’s predictions and ground truth distributions.
In general, our findings highlight a clear difficulty across models in accurately handling datasets that require complex, sparse, or timely executed actions, particularly when special non-movement actions are involved. The macro recall performance appears directly correlated to how closely model predictions distribution aligns with the ground truth action distribution.
IV-B2 Action-Class Specific Performance Across Models
A detailed analysis of recall performance by individual action classes offers valuable insights into model-specific strengths, weaknesses, and intrinsic biases in predicting discrete actions. As seen in Figure 10 action classes 4 through 10 generally displayed higher recall scores and greater variance among models, indicating varying degrees of predictive confidence. In contrast, action classes 0 to 3 and 11 to 12 consistently yielded low recall values across all models, demonstrating limited predictive capability.
Figure 10: Class-wise recall averaged across all subdatasets for all 5 models. Recall and Variance on classes 4-10 is relatively higher, and is low on classes 0-3, and 11-12.
GPT-4o exhibited notable recall performance on action class 7 (”Right” movement), achieving a high recall of approximately 0.22. This suggests a directional bias likely influenced by the prevalent rightward progression in many Procgen tasks. Conversely, GPT-4o struggled significantly with actions 4, 10, 11, and 12, displaying near-zero recall, highlighting clear limitations in handling certain peripheral or specialized actions.
OpenVLA predominantly favored action class 4 (”Do Nothing”), recording the highest recall ( 0.32). This bias arises from a combined influence of its decoding mechanism and Procgen action space sizes. Most Procgen datasets have an action space of size 9 or 10, and due to unnormalization of OpenVLA’s predicted actions using the dataset statistics, 4 can correspond to 0 in the normalized range which is the center of OpenVLA’s output distribution. However, OpenVLA encountered considerable difficulty predicting peripheral classes (0, 1, 8, 10, and 11), with recall scores as low as 0.03, indicating a substantial predictive bias toward central actions.
Pi0 Base similarly gravitated toward central action classes (4, 5, 6, and 7) due to its normalization-unormalization decoding approach, resulting in moderate recall values (  0.12 - 0.16). Its recall sharply declined for peripheral and special action classes (0, 1, 2, and 12), dropping to approximately 0.02. Compared to OpenVLA, Pi0 Base’s diffusion-based approach led to less concentrated predictions, distributing recall more evenly but still inadequately across peripheral classes.
Pi0 FAST displayed a pronounced bias towards higher-numbered action classes (8, 9, and 10), with recall scores around 0.10. This bias reflects its difficulty generalizing to OOD data, often producing higher normalized action predictions. Conversely, Pi0 FAST severely underperformed on lower-numbered and mid-range actions (0, 2, 3, 7, and 11), frequently achieving near-zero recall, emphasizing its limitations in balanced action class predictions.
GPT-4.1 demonstrated a distinct bias toward action class 7 (”Right” movement), achieving exceptionally high recall ( 0.40), consistent with the directional tendencies of many Procgen environments. It also showed moderate recall ( 0.20) for special action class 9. However, GPT-4.1 struggled notably with several other action classes, particularly actions 0, 1, 4, 10, and 11, consistently resulting in minimal recall ( 0.01).
These results clearly indicate model-specific biases towards particular action classes, underscoring the importance of targeted adjustments in model training and decoding methods to foster more balanced and generalized performance across the entire action space.
IV-B3 Models’ Prediction Collapse to Few Action Classes
A consistent observation across all evaluated models was their tendency to collapse predictions predominantly into a limited subset of action classes. Analysis of union confusion matrices as seen in Figure 11 aggregated over all 16 Procgen subdatasets highlighted clear default prediction behaviors.
Specifically, each model exhibited distinct biases toward particular default actions:
Lists are easy to create:
  • OpenVLA frequently defaulted to predicting action classes 4 (”Do Nothing”) and 6 (”RIGHT + DOWN”).
  • Pi0 Base predominantly predicted action classes clustered between 4 to 7 (”Do Nothing,” ”UP,” ”RIGHT + DOWN,” ”RIGHT”) and occasionally between 5 to 8 (”UP,” ”RIGHT + DOWN,” ”RIGHT,” ”RIGHT + UP”).
  • Pi0 FAST consistently favored action classes 6 (”RIGHT + DOWN”) and 8 (”RIGHT + UP”).
  • GPT-4o exhibited a strong bias towards action class 7 (”RIGHT”).
  • GPT-4.1 frequently defaulted to action classes 7 (”RIGHT”) and 2 (”LEFT + UP”).
Figure 11: Union confusion matrices depicting frequency of class predictions and ground truth matches. Order from top to bottom - GPT 4o, OpenVLA, Pi0 Base, Pi0 FAST, GPT 4.1
Notably, as seen in Figure 12, across all models and subdatasets, action class 4 (”Do Nothing”) exhibited significantly higher precision compared to other classes. This elevated precision for class 4 likely originates from its role as a default assignment for expert actions that fall outside the environment-specific action spaces when creating ground truth labels, thus inflating its representation in datasets.
Figure 12: Class-wise precision averaged across all subdatasets for all 5 models. Precision is significantly higher for 4 across all models
For autoregressive models like GPT-4o and GPT-4.1, the frequent default to action class 4 may also reflect insufficient understanding of the prompt and task, causing uncertainty in predictions and defaulting to ”Do Nothing.” Additionally, the observed strong preference for ”RIGHT” (action class 7) aligns with the prevalent directional progression in many Procgen environments.
In contrast, OpenVLA, Pi0 Base, and Pi0 FAST primarily defaulted towards actions situated near the center of their normalized prediction range (typically zero). Upon unnormalization, these central predictions commonly mapped to action class 4 due to rounding effects and the typical size (<<<= 10 actions) of Procgen datasets.
These findings underscore critical biases inherent in model designs and highlight opportunities for refining training approaches and decoding methods to achieve balanced action distributions and improve generalization to unseen action spaces.
IV-C Models’ Predictions Collapse on Specific Action Classes with OOD Data
IV-C1 Patterns of Collapse Influenced by Decoding Techniques
The behavior of models on out-of-distribution (OOD) data revealed clear patterns of prediction collapse, significantly influenced by their underlying decoding strategies. We specifically examined how diffusion sampling and autoregressive (AR) decoding methods shaped these patterns.
Diffusion models inherently differ from AR models due to their generative approach, which aims to reconstruct data from noise by modeling the entire distribution of training data [34]. In multi-class classification tasks with mutually exclusive single-correct answers, diffusion models respond distinctly when encountering OOD inputs. Primarily, diffusion models excel at reconstructing data similar to their training distribution. However, upon encountering OOD inputs, their reconstruction accuracy deteriorates markedly due to difficulty in mapping the unfamiliar inputs onto the learned data manifold. Consequently, predictions on OOD data are characterized by uncertainty, with predictions distributed across multiple classes instead of sharply peaked at a single class. Visual inspection of confusion matrices from Pi0 Base on datasets like Bigfish, Climber, and Maze as seen in Figures 24, 25, and 26, supports this finding, demonstrating distributed prediction frequencies rather than dominance by any one class.
In contrast, AR models sequentially predict outputs conditioned on previous outputs, typically applying softmax at the final step for classification tasks. A known limitation of this approach is an inherent tendency toward overconfidence in OOD scenarios. Despite inputs being significantly divergent from training data, AR models frequently produce sharply peaked probability distributions, arbitrarily favoring one class. This behavior arises from their inductive biases, where the softmax activation overestimates probabilities for certain classes. This issue is evidenced clearly in the union confusion matrices seen in Figure 11. Each AR model consistently collapses its predictions onto particular classes, even with nonsensical or significantly divergent inputs, as observed in the individual per-subdataset confusion matrices in the Appendix section VII-A for OpenVLA’s outcomes on Chaser and Jumper, Pi0 FAST’s results on Bigfish, Chaser, and Heist, and GPT 4x’s performance on Heist, Maze, Coinrun and Miner.
The prediction collapse behavior also varies significantly due to differences in tokenization methods employed by AR models. Specifically, OpenVLA and Pi0 FAST differ fundamentally in their tokenization and decoding strategies, influencing their OOD performance distinctly.
Pi0 FAST utilizes the Discrete Cosine Transform (DCT) [1] in combination with Byte Pair Encoding (BPE) [64], focusing on low-frequency components that represent overarching shape of the signals. This frequency-domain approach, particularly the strategy of flattening DCT coefficients prioritizing low-frequency information, results in concentrated predictions due to stronger global pattern capture and robust representation against input noise. This approach enhances consistency and stability in action predictions, as evidenced by concentrated class predictions observed in the confusion matrices.
Conversely, OpenVLA employs a discrete binning approach, mapping continuous action spaces to discrete bins directly, thus not capturing global patterns in any sense when evaluated in a zero-shot setting. Each action dimension or timestep is processed independently, which leads to a more diffused spread of predictions. OpenVLA exhibits sensitivity to variations and input noise, resulting in more evenly distributed probability mass across action classes, as reflected in the broader spread observed in its confusion matrices.
Overall, our analysis highlights how decoding methods and tokenization strategies critically impact models’ behaviors in OOD scenarios, indicating the need for targeted adjustments depending on the specific application requirements for generalization performance.
V Generalization Barriers in Vision-Language-Action Models: Dataset, Architecture, and Processing Techniques
V-1 Differences Between Training Data and Procgen
Dataset Domain Discrepancies
A fundamental obstacle limiting model performance arises from substantial differences between the training datasets and the Procgen evaluation dataset. Procgen environments are procedurally generated, Atari-like 2D games designed to evaluate visual and motor skills of reinforcement learning (RL) agents. Each subdataset within Procgen exhibits diverse environment layouts, tasks, objectives, reward structures, and discrete action spaces, predominantly involving directional movements and specific game-related interactions.
Conversely, the training datasets vary significantly across models:
GPT 4x models were primarily trained on expansive web-scale vision-language data, lacking a dedicated focus on action-oriented interactions.
OpenVLA is a vision-language model specifically fine-tuned on robotics data from the Open X-Embodiment dataset, comprising approximately 970,000 robot demonstrations. These demonstrations are strictly limited to continuous robotic manipulation tasks recorded from third-person camera views.
Pi0 models also leverage robotics trajectory data featuring continuous action spaces. Their training incorporates a subset of OpenX-Embodiment alongside a private dataset containing approximately 903 million timesteps, significantly sourced from single and dual-arm robotic manipulation scenarios.
The substantial domain gap—real-world robotic manipulation versus simulated game interaction, continuous versus discrete action spaces, and physical manipulation versus simple directional controls—makes zero-shot adaptation particularly challenging for vision-language-action (VLA) models.
Action Space Awareness
Each model exhibits varying degrees of awareness regarding the action space constraints:
GPT 4x models are explicitly informed through prompts about valid action values and corresponding verbal descriptions.
OpenVLA employs output clamping, effectively limiting predictions strictly within valid action ranges upon unnormalization based on dataset statistics.
Pi0 Base and Pi0 Fast rely on training-induced normalization but lack explicit clamping, occasionally resulting in invalid predictions outside permissible action ranges.
Proprioceptive State Handling
Procgen datasets lack proprioceptive state information, a critical input for Pi0 models during training. To compensate, zero arrays replace missing proprioceptive states during inference, negatively impacting Pi0 model performance by introducing unnatural, non-representative states. While OpenVLA optionally accepts proprioceptive states, the absence of these in Procgen data inherently limits the richness of information available for inference.
Image Input Variability
Significant discrepancies in image inputs further exacerbate model adaptation difficulties:
Resolution mismatch: Training images for OpenVLA and Pi0 models are at 224x224 resolution, whereas Procgen images are limited to 64x64 pixels, necessitating zero-padding and resizing, likely impairing visual perception and subsequent performance.
View differences: OpenVLA expects a single image input, matching the Procgen input structure, potentially enhancing its adaptability. In contrast, Pi0 models trained on multiple views suffer due to missing additional viewpoints, where zero arrays substitute these inputs, further degrading performance.
V-2 Difficulty Adapting to Increased Image Complexity
We quantified image complexity through Shannon entropy [66] and Delentropy [57] across the first 20 episodes per dataset as seen in Figures 13 and 14, revealing distinctive performance patterns. The formula used to calculate the entropies can be seen in Appendix section VII-B:
Figure 13: Entropy measures VS Model Macro Recall
Figure 14: Correlation matrix for Entropy measures VS Model Macro Recall
GPT-4 models (4o and 4.1) demonstrated moderate to weak negative correlations (Shannon entropy: -0.41 and -0.28; Delentropy: -0.351 and -0.25, respectively), suggesting poorer performance with increasing complexity.
OpenVLA and Pi0 Base exhibited negligible correlations with entropy metrics (OpenVLA: 0.071 Shannon, 0.17 Delentropy; Pi0 Base: 0.047 Shannon, 0.063 Delentropy), implying relative independence from image complexity.
Pi0 Fast uniquely displayed moderate positive correlations (Shannon and Delentropy: 0.44), indicating improved performance with higher image complexity.
These findings highlight that image complexity impacts GPT-4 models negatively, benefits Pi0 Fast, and minimally affects OpenVLA and Pi0 Base. The moderate correlation strength indicates complexity alone does not fully explain model performance variability.
V-3 Impact of VLA Training, Decoding, and Output Processing Techniques
Distinct training, decoding, and output processing strategies significantly shape model generalization capabilities:
OpenVLA utilizes a robust output processing approach, training to output actions within a normalized range of [-1,1], with strict clamping enforced. Consequently, its predictions, upon unnormalization, consistently map within valid action spaces, eliminating invalid outputs and markedly enhancing zero-shot generalization.
Pi0 Base also targets normalized outputs within [-1,1] without enforced clamping. Its diffusion-based architecture inherently predicts actions ”close” to familiar training distributions. This approach yields relatively fewer invalid outputs and moderately effective generalization compared to purely autoregressive models but remains inferior to OpenVLA.
Pi0 FAST, while similarly trained on normalized action outputs, employs autoregressive decoding without clamping, lacking the intrinsic action proximity measure found in diffusion models. Consequently, Pi0 FAST predictions deviate significantly from valid actions when generalization fails, resulting in extremely high invalid output rates and poor overall performance, except for moderate macro precision.
Collectively, these findings underscore critical architectural and training considerations necessary for enhancing generalization across diverse and unseen discrete action environments.
VI Conclusion
In this study, we systematically evaluated the generalization capabilities of contemporary VLAs and VLMs across procedurally generated discrete-action environments from the Procgen dataset in a zero-shot setting. Our analysis highlighted significant limitations arising from architectural constraints, training paradigms, and input-output biases inherent to the models. The stark domain discrepancy between training data—primarily continuous-action robotics datasets, and general web-scale vision language data—and discrete-action game environments emerged as a critical barrier to effective zero-shot generalization.
We identified notable differences in model behaviors linked directly to their architectures, training strategies, and input/output processing techniques. OpenVLA’s robust action-space clamping technique consistently provided superior generalization, minimizing invalid outputs and exhibiting relative resilience to out-of-distribution scenarios. Conversely, autoregressive models like GPT-4x displayed substantial difficulties in generalizing, especially under complex image conditions, and frequently defaulted to overly simplistic or biased action choices. Additionally, Pi0 models showed intermediate performance influenced heavily by their diffusion-based (Pi0 Base) or autoregressive (Pi0 FAST) decoding methods, with Pi0 FAST notably sensitive to image complexity and unable to restrict majority of its predictions to a desired output range.
Our findings underscore the necessity for architectural innovations, refined training methodologies, and enhanced output processing techniques to bridge the gap between diverse action domains. Future research should prioritize developing more generalized training datasets that better reflect the variety of potential application environments, alongside methods to adaptively handle different forms of action representations. These advancements hold promise for enabling VLAs to operate effectively across an increasingly diverse and unpredictable range of real-world tasks.